# How Does the Textual Information Affect the Retrieval of Multimodal In-Context Learning?

### v1

@@ 4.1 Datasets
Following Alayrac et al. (2022), we focus on three representative multimodal tasks and the details about the datasets used for these tasks as follows. MS COCO Lin et al. (2015) for image captioning, OK-VQA Marino et al. (2019) for Visual Question Answering, and HatefulMemes Kiela et al. (2021) for rank classification. Accuracy on the test split is measured for OK-VQA. In MSCOCO, evaluation utilizes CIDEr scores Vedantam et al. (2015) on the Karpathy-test split. For HatefulMemes, the AUC ROC is calculated. All results are averaged with three runs. For further details of these downstream tasks and corresponding datasets, please refer to Appendix A.3 A.4.
4.2 Multimodal Large Language Models
The experiments conducted herein utilize the OpenFlamingo-3B framework Awadalla et al. (2023). The architecture of OpenFlamingo encompasses a fixed large language model featuring a decoder-only configuration (for instance, MPT [31]), accompanied by a static visual encoder (such as CLIP-ViT Radford et al. (2021)), which is succeeded by a trainable perceiver resampler. To facilitate the integration of visual and linguistic data, trainable cross-attention layers are strategically interspersed among the pre-trained language model layers. OpenFlamingo underwent pre-training on the LAION-2B Schuhmann et al. (2022) and Multimodal C4 Zhu et al. (2023) datasets and its performance is on par with that of Flamingo Alayrac et al. (2022), demonstrating its competitive edge. The justification for selecting OpenFlamingo as the foundational model is provided in Appendix A.2.
@@ 4.3 Compared Methods
The analysis evaluates various methodologies, including the Random baseline, RICES unsupervised retrieval (Q-I-M-I), Q-T-M-T textual variant, and two proposed approaches: Multimodal Unsupervised In-context Examples Retrieval (MUIER) using all multimodality aspects, and Multimodal Supervised In-context Examples Retrieval (MSIER) enhancing CLIP’s dual encoders. MSIER undergoes 30 epochs of training with the AdamW optimizer, starting at a learning rate 1e-5, subject to reduction using the cosine annealing rule. For further details of these methods, please refer to Appendix A.1.
@@ 4.4 Main Results
Compared with Randon and RICES, The Multimodal Unsupervised In-context Examples Retrieval (MUIER) method shows further enhancements, underscoring the benefits of integrating images and text in a retrieval strategy without explicit supervision. This approach enables the model to better understand in-context examples, with the textual content offering additional direction, thus improving the retrieval of more relevant examples. The Multimodal Supervised In-context Examples Retrieval (MSIER) method outperforms all with a significant rise in performance, evidenced by a $5.52$ increase in the CIDEr score of MS COCO dataset. This suggests that a refined retrieval mechanism, informed by the foundational knowledge of multimodal large language models (MLLMs) and tailored to select correct image-text pairings, yields superior learning outcomes.
@@ 4.5 Further Analysis
Figure 6: Multimodal in-context examples retrieved by RICES, MUIER and MSIER. In each grid, the first four blue rows contain the prompt while the orange row contains the query image and prediction. Further outcomes are detailed in the Appendix A.5
How does the MSIER improve M-ICL specifically? To address this query, we employ a multimodal analysis of the in-context examples identified by MUIER and MSIER as depicted in Figure 6. Our examination concentrates on image captioning by selecting one example from the MS COCO dataset where the number of shots is 4. Concretely, the columns from left to right delineate the retrieved in-context example with RICES, MUIER, and MSIER separately. The blue rows are the retrieved multimodal in-context examples, that is, a pair comprising input and output, whereas the subsequent orange rows present the query image alongside the model’s prediction. Through a comparative analysis of the in-context examples selected by RICES, MUIER, and MSIER, we elucidate the superior performance of MSIER over MUIER; the examples identified by MSIER exhibit a closer semantic similarity to the queries.
How does the order of in-context examples affect M-ICL? To understand if changing the order of multimodal in-context examples makes a difference, we fix the number of in-context examples to 3, and evaluate all possible permutations. As shown in Figure 5, the standard deviation is generally small, so the order is not a concern as long as high-quality examples are chosen.
Transferability The compositional characteristics of natural language and images are general, meaning the retriever may exploit similar knowledge in different tasks or scoring MLLMs. This motivates us to explore whether the proposed MSIER trained on one dataset or MLLM scorer can be directly transferred to others without further tuning. This is a practical research question as training a retriever for each dataset or MLLM scorer can be costly in real applications.
  • Datasets To measure the transferability of supervised retrievers between datasets, we evaluate the performance of MSIER trained on OK-VQA and HatefulMemes datasets on MS COCO dataset with different numbers of shots of retrieved in-context examples. The result in Table 3 manifests some transferability of our proposed MSIER between multiple multimodal datasets. Enhanced performance of MSIER with training on the OK-VQA dataset suggests that data form and volume significantly influence MSIER’s effectiveness.
@@
Table 4: Results of M-ICL performance of random selection, RICES, and MUIER-9B and MSIER-3B method on MS COCO dataset with OpenFlamingo-9B as scorer.
@@ 4.6 Ablation Study
This section presents an extensive ablation study based on the OKVQA and MS COCO dataset to verify the best setting for our proposed model’s components,
Impact of Number of Candidates We proceed to assess the performance of MSIER across a range of candidate numbers. In MSIER, we pass the candidates to a scoring LM and label the top-$K$ and the bottom-$K$ as positive and negative examples respectively. If the number of candidates is not large enough, it will not hold sufficient information for contrastive learning; conversely, if we set the bank with a large size, it will contain a quantity of irrelevant items. The result presents $K$=5 as the best choice on OK-VQA dataset.
@@
Impact of Textual Information in Retrieved In-context Examples The role of textual information during the evaluation of M-ICL remains unexplored for our proposed MSIER. We find that compared to the standard scenario, the replacement of texts in the in-context examples hugely impacts the M-ICL performance. To be specific, employing a mask on half of the captions through the random selection method results in a significant decrease in the CIDEr score. Correspondingly, the side effect of the removal of texts decreases from MUIER method to MSIER method compared with random retrieval.
@@
Impact of Backbone Retriever To understand if using a different backbone would make a difference, we further evaluate our retrieval methods, MUIER and MSIER, on the image captioning benchmark using another backbone: ALIGN Jia et al. (2021) pre-trained using contrastive learning on noisy image-text pairs. The results are reported in Table 7, where the MUIER and MSIER methods do not demonstrate explicit advantages over random selection, suggesting that the ALIGN model does not offer retrieval benefits for the M-ICL of the OpenFlamingo model. This is due to OpenFlamingo’s use of the CLIP model as its vision encoder, resulting in poor contextual information from examples retrieved primarily by the ALIGN model, leading to little enhancement in M-ICL performance.
@@
Random 77.02 84.13 88.96 92.32 85.61
MUIER (Q-I-M-IT) 78.78 84.1 89.22 92.07 86.04
MSIER (Q-I-M-IT) 78.17 83.79 89.38 94.02 86.34
@@ 5 Conclusion
In this study, we conduct an extensive evaluation of textual information’s role in in-context example retrieval for multimodal in-context learning. We introduce MUIER and MSIER methodologies that incorporate textual data in both unsupervised and supervised retrieval processes. Our experiments across three multimodal tasks demonstrate that integrating text significantly enhances the efficiency of example selection, yielding substantial improvements over existing approaches. Future research could delve into optimizing inter-modal interactions, leveraging our findings for more effective multimodal in-context learning retrieval strategies.
6 Limitation
@@ Acknowledgements
Appendix A Appendix
@@ A.1 Compared Methods
  • The novel approach introduced by this research, termed Multimodal Unsupervised In-context Examples Retrieval (MUIER), employs readily accessible features to identify the most similar examples via search, based on all multimodality aspects.
  • Multimodal Supervised In-context Examples Retrieval (MSIER) represents our secondary proposal, which involves the enhancement of CLIP’s dual encoders. This enhancement process is facilitated through direct optimization to improve M-ICL performance. The supervised model undergoes training for 30 epochs utilizing the AdamW optimizer. The initial learning rate is established at 0.00001, subject to reduction according to the cosine annealing rule.
A.2 MLLM backbone
@@ A.4 Datasets
  • MSCOCO A large-scale dataset for object detection, segmentation, and captioning. It provides diverse images with complex scenes and multiple objects in context. Annotations include object segmentation, recognition, and image captions. The dataset contains approximately 330K images, with 1.5 million labeled instances across 80 object categories. We utilize Karparthy’s split: training (83K images), validation (5K images), and test (5K images) sets.
  • OK-VQA A dataset designed for open-ended Visual Question Answering that requires external knowledge beyond image content. It features questions demanding multi-modal knowledge, combining visual cues with general world knowledge. The dataset comprises over 14,000 images sourced from the MSCOCO dataset, with around 14,000 questions spanning 10 categories. We leverage its structured format, which includes a balanced mix of 9,009 training and 5,046 testing questions, to assess the capability of models to integrate visual understanding with external knowledge sources.
  • HatefulMemes A dataset designed for the detection and classification of hate speech in multimodal content. It uniquely combines textual and visual elements to challenge models in understanding complex, nuanced expressions of hate speech. The dataset consists of 10,000+ meme images, annotated for hate speech detection, with a distribution of 8,500 for training and 3,000 for testing. The Hateful Memes dataset provides a significant challenge in discerning subtle contextual cues and cultural references, requiring advanced multimodal analysis capabilities.
A.5 More Evaluation Examples

### v2

@@ 4.1 Datasets
Following Alayrac et al. (2022), we focus on three representative multimodal tasks and the details about the datasets used for these tasks as follows. MS COCO Lin et al. (2015) for image captioning, OK-VQA Marino et al. (2019) for Visual Question Answering, and HatefulMemes Kiela et al. (2021) for rank classification. Accuracy on the test split is measured for OK-VQA. In MSCOCO, evaluation utilizes CIDEr scores Vedantam et al. (2015) on the Karpathy-test split. For HatefulMemes, the AUC ROC is calculated. All results are averaged with three runs. For further details of these downstream tasks and corresponding datasets, please refer to Appendix A.3 and A.4. The experiments conducted herein utilize the OpenFlamingo-3B framework Awadalla et al. (2023). Comprehensive details regarding OpenFlamingo can be found in Appendix A.2.
@@ 4.2 Compared Methods
The analysis evaluates various methodologies, including the Random baseline, RICES (Q-I-M-I), Q-T-M-T textual variant, Multimodal Unsupervised In-context Examples Retrieval (MUIER) using all multimodality aspects, and the proposed approach: Multimodal Supervised In-context Examples Retrieval (MSIER) enhancing CLIP’s dual encoders. MSIER undergoes 30 epochs of training with the AdamW optimizer Kingma and Ba (2017), warming up to a peak learning rate 1e-5, subject to reduction using the cosine annealing rule. For further details of these methods, please refer to Appendix A.1.
@@ 4.3 Main Results
As shown in Table 2, compared with Random and RICES, MUIER method shows further enhancements, underscoring the benefits of integrating images and text in a retrieval strategy without explicit supervision. This approach enables the model to better understand in-context examples, with the textual content offering additional direction, thus improving the retrieval of more relevant examples. MSIER method outperforms all with a significant rise in performance, evidenced by a $5.52$ increase in the CIDEr score of MS COCO dataset. This suggests that a refined retrieval mechanism, informed by the foundational knowledge of multimodal large language models (MLLMs) and tailored to select correct image-text pairings, yields superior learning outcomes. Furthermore, MSIER with only 4 shots outperforms random selection with 32 shots. Thus, employing our method can enable MLLMs to achieve comparable M-ICL performance with limited memory usage.
@@ 4.4 Further Analysis
Figure 6: Multimodal in-context examples retrieved by RICES, MUIER and MSIER. In each grid, the first four blue rows contain the prompt while the orange row contains the query image and prediction. Further outcomes are detailed in the Appendix B.
How does the MSIER improve M-ICL specifically? To address this query, we employed a multimodal analysis of the in-context examples identified by MUIER and MSIER as depicted in Figure 6. Our examination concentrates on image captioning by selecting one example from the MS COCO dataset where the number of shots is 4. Concretely, the columns from left to right delineate the retrieved in-context example with RICES, MUIER, and MSIER separately. The blue rows are the retrieved multimodal in-context examples, that is, a pair comprising input and output, whereas the subsequent orange rows present the query image alongside the model’s prediction. In the given figure, examples identified by MSIER exhibit a closer semantic similarity to the queries than RICES and MUIER. Consequently, the MLLM made a more accurate prediction that includes the caption "a chocolate covered peanut", which is not captured by previous predictions using RICES and MUIER methods for retrieving in-context examples.
How does the order of in-context examples affect M-ICL? To understand if changing the order of multimodal in-context examples makes a difference, we fixed the number of in-context examples to 3, and evaluate all possible permutations. As shown in Figure 5, the standard deviation is generally small, so the order is not a concern as long as high-quality examples are chosen.
How is the transferability of proposed MSIER? The compositional characteristics of natural language and images are general, meaning the retriever may exploit similar knowledge in different tasks or scoring MLLMs. This motivates us to explore whether the proposed MSIER trained on one dataset or MLLM scorer can be directly transferred to others without further tuning. This is a practical research question as training a retriever for each dataset or MLLM scorer can be costly in real applications.
  • Datasets To measure the transferability of supervised retrievers between datasets, we evaluated the performance of MSIER trained on OK-VQA and HatefulMemes datasets on MS COCO dataset with different numbers of shots of retrieved in-context examples. The result in Table 3 manifests some transferability of our proposed MSIER between multiple multimodal datasets. Enhanced performance of MSIER with training on the OK-VQA dataset suggests that data form and volume significantly influence MSIER’s effectiveness.
@@
Table 4: Results of M-ICL performance of random selection, RICES, and MUIER-9B and MSIER-3B method on MS COCO dataset using OpenFlamingo-9B for inference.
@@ 4.5 Ablation Study
This section presents an extensive ablation study based on the OKVQA and MS COCO datasets to verify the best setting for our proposed model’s components.
Impact of Different Modality Encoders in CLIP Unlike ICL retrieval in NLP and CV, M-ICL retrieval employs multiple encoders in its backbone retriever to extract features from different modalities. To investigate the roles of the text and image encoders within the base CLIP model in our proposed MSIER framework, we selectively freeze one encoder during MSIER training. Table 5 demonstrates that the update on image encoder has a more significant effect on M-ICL retrieval performance improvement because freezing the image encoder results in a notable performance drop compared to the original MSIER (1.39 $\downarrow$). In contrast, freezing only the text encoder yields similar or even improved retrieval performance, suggesting that the original trained MSIER may have overfitted to some degree on the text encoder.
Method 4 shots 8 shots 16 shots 32 shots Avg
MUIER (CLIP) 92.78 98.93 103.65 107.24 100.65
MSIER (CLIP) 100.58 105.09 108.41 110.58 106.17
MSIER (CLIP-Freeze T) 100.82 105.90 109.17 110.24 106.53
MSIER (CLIP-Freeze I) 98.13 103.37 107.59 110.01 104.78
Table 5: Comparison of M-ICL performance of MSIER method on MS COCO dataset with freezing different modality encoders of the backbone retriever.
Impact of Number of Candidates We proceed to assess the performance of MSIER across a range of candidate numbers. In MSIER, we passed the candidates to a scoring LM and label the top-$K$ and the bottom-$K$ as positive and negative examples respectively. If the number of candidates is not large enough, it will not hold sufficient information for contrastive learning; conversely, if we set the bank with a large size, it will contain a quantity of irrelevant items. The results in Table 6 present $K$=5 as the best choice on OK-VQA dataset.
@@
Impact of Textual Information in Retrieved In-context Examples The role of textual information during the evaluation of M-ICL remains unexplored for our proposed MSIER. In Table 7, we find that compared to the standard scenario, the replacement of texts in the in-context examples hugely impacts the M-ICL performance. To be specific, employing a mask on half of the captions through the random selection method results in a significant decrease in the CIDEr score. Correspondingly, the side effect of the removal of texts decreases from MUIER method to MSIER method compared with random retrieval.
@@
Impact of Backbone Retriever To investigate whether using a different backbone model could improve performance, we further evaluated retrieval methods, MUIER and MSIER, on the image captioning benchmark by utilizing an alternative backbone: EVA-02-CLIP Sun et al. (2023), a model enhances the original CLIP model with advanced pre-training techniques. The results, reported in Table 8, show that MUIER and MSIER methods employing EVA-02-CLIP(-L/14) as the backbone retriever demonstrated better performance. This suggests that the EVA-02-CLIP model offers retrieval benefits for the M-ICL of the OpenFlamingo model, with improved feature extraction for both textual and visual information. Furthermore, the proposed MSIER method exhibited a corresponding performance improvement when using the enhanced retriever, explicitly validating MSIER’s excellent transferability and ability to achieve better results as multimodal encoders advance. Additional experiments can be found in Appendix 9.
@@
MUIER (CLIP) 92.78 98.93 103.65 107.24 100.65
MSIER (CLIP) 100.58 105.09 108.41 110.58 106.17
MUIER (EVA-02-CLIP) 94.02 101.70 107.06 109.71 103.12
MSIER (EVA-02-CLIP) 104.20 107.44 111.61 113.07 109.08
Table 8: Comparison of M-ICL performance of MUIER and MSIER method on MS COCO dataset with EVA-CLIP2 model as the backbone retriever.
@@ 5 Conclusion
In this study, we conducted an extensive evaluation of textual information’s role in unsupervised and supervised in-context example retrieval for multimodal in-context learning. We introduce a novel MSIER methodology that incorporates MLLMs’ self-contained information to train a supervised prompt retriever. Our experiments across three multimodal tasks demonstrate that integrating text modality and the foundational knowledge of MLLMs significantly enhances the efficiency of example selection, yielding substantial improvements over existing approaches. Moreover, MSIER demonstrates high efficiency by outperforming benchmarks with far fewer examples, while also exhibiting strong transferability by remaining effective for larger models after training on smaller ones, thereby improving cost-efficiency for large-scale inference. Future research could delve into optimizing inter-modal interactions, leveraging our findings for more effective multimodal in-context learning retrieval strategies.
6 Limitations
@@ Acknowledgements
Appendix A More Experimental Details
@@ A.1 Compared Methods
  • The standard unsupervised approach used in this research, termed Multimodal Unsupervised In-context Examples Retrieval (MUIER), employs readily accessible features to identify the most similar examples via search, based on all multimodality aspects.
  • Multimodal Supervised In-context Examples Retrieval (MSIER) represents our secondary proposal, which involves the enhancement of CLIP’s dual encoders. CLIP ViT-L-14 is applied for all experiments. This enhancement process is facilitated through direct optimization to improve M-ICL performance. The supervised model undergoes training for 30 epochs utilizing the AdamW optimizer. The initial learning rate is established at 0.00001, subject to reduction according to the cosine annealing rule.
A.2 Multimodal Large Language Model
The architecture of OpenFlamingo encompasses a fixed large language model featuring a decoder-only configuration (for instance, MPT [31]), accompanied by a static visual encoder (such as CLIP-ViT Radford et al. (2021)), which is succeeded by a trainable perceiver resampler. To facilitate the integration of visual and linguistic data, trainable cross-attention layers are strategically interspersed among the pre-trained language model layers. OpenFlamingo underwent pre-training on the LAION-2B Schuhmann et al. (2022) and Multimodal C4 Zhu et al. (2023) datasets and its performance is on par with that of Flamingo Alayrac et al. (2022), demonstrating its competitive edge.
@@ A.4 Datasets
  • MSCOCO is a large-scale dataset for object detection, segmentation, and captioning. It provides diverse images with complex scenes and multiple objects in context. Annotations include object segmentation, recognition, and image captions. The dataset contains approximately 330K images, with 1.5 million labeled instances across 80 object categories. We utilize Karparthy’s split: training (83K images), validation (5K images), and test (5K images) sets.
  • OK-VQA is a dataset designed for open-ended Visual Question Answering that requires external knowledge beyond image content. It features questions demanding multi-modal knowledge, combining visual cues with general world knowledge. The dataset comprises over 14,000 images sourced from the MSCOCO dataset, with around 14,000 questions spanning 10 categories. We leverage its structured format, which includes a balanced mix of 9,009 training and 5,046 testing questions, to assess the capability of models to integrate visual understanding with external knowledge sources.
  • HatefulMemes A dataset constructed for the detection and classification of hate speech in multimodal content. It uniquely combines textual and visual elements to challenge models in understanding complex, nuanced expressions of hate speech. The dataset consists of 10,000+ meme images, annotated for hate speech detection, with a distribution of 8,500 for training and 3,000 for testing. The Hateful Memes dataset provides a significant challenge in discerning subtle contextual cues and cultural references, requiring advanced multimodal analysis capabilities.
A.5 Retrieval Methods
In our main experiments, MUIER and MSIER utilized all modalities to calculate similarity scores and directly selected the most similar items as in-context examples. Recently, some newly proposed retrieval methods have emerged, such as Mixed Modality In-Context Example Selection Chen et al. (2023a), which uses visual features to retrieve a set of top-N candidate examples and then employs textual information to retrieve the final in-context examples (with varying numbers of examples, such as 4, 8, 16, or 32). We also validated the efficiency of MSIER in this scenario, and Table 9 shows that, although we did not train a new MSIER specifically designed for the new retrieval method, it still demonstrated better performance on the VQA task. This further highlights the generality of our proposed MSIER framework.
Method 4 shots 8 shots 16 shots 32 shots Avg
MMICES-CLIP 32.96 33.60 35.89 36.13 34.65
MMICES-MSIER 33.08 34.36 35.92 36.84 35.05
Table 9: Comparison of M-ICL performance of MMICES with CLIP as retriever, and MMICES with MSIER as retriever on MS COCO dataset. We chose the top 50 candidates for further selection of in-context examples in this experiment.
Impact of Backbone Retriever We further evaluated the retrieval methods MUIER and MSIER using a different backbone: ALIGN Jia et al. (2021), a model pre-trained on noisy image-text pairs using contrastive learning. The results, reported in Table 10, show that the MUIER and MSIER methods did not demonstrate explicit advantages over random selection. This suggests that the ALIGN model does not offer retrieval benefits for the M-ICL of the OpenFlamingo model. This can be attributed to the fact that OpenFlamingo utilizes the CLIP model as its vision encoder, resulting in poor contextual information from examples primarily retrieved by the ALIGN model, leading to little enhancement in M-ICL performance.
Method 4 shots 8 shots 16 shots 32 shots Avg
Random 77.02 84.13 88.96 92.32 85.61
MUIER (ALIGN) 78.78 84.10 89.22 92.07 86.04
MSIER (ALIGN) 78.17 83.79 89.38 94.02 86.34
Appendix B More Evaluation Examples